# Efficient Multi-modal Large Language Models via Visual Token Grouping

### v1

@@ 2.2 Efficient inference for LLM/MLLM
The auto-regressive nature of LLMs poses a great challenge to the deployment of LLMs. The quadratic complexity of computing the attention makes the generation process much slower when the input token is longer. StreamingLLM [25] and FastGen [9] prune the redundant attention computation to simplify the computation. Despite their success, they are designed for the single-modal LLM and have not proven successful when it comes to scenarios involving tokens from other modalities. For improving the efficiency of MLLMs, various works [11, 34, 32] adopt Q-Former to map the images to fixed-length tokens. In the meantime, many works try to train smaller MLLMs with smaller backbone [30, 33] to handle the scenario with less computational resources, MoE-LLaVA [14] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance. Another line of research tries to reduce the number of visual tokens while keeping the backbone of LLM unchanged. Since LLMs contribute most to the computation, the number of input tokens to the LLM becomes the bottleneck of the inference cost. Recently, DeCo [28] employs 2D adaptive average pooling to down-sample the visual tokens at the patch level. Honeybee [4] proposes to use ResNet Block and deformable attention to conduct the abstraction of the vision tokens. VoCo-LLaMA [29] compress the vision tokens using the LLMs and leverage the attention distillation to let LLMs restore information from the specially defined VoCo tokens instead of the whole image tokens.
@@ 3.1 Recap of Multi-modal Large Language Models
[…]
@@ 3.2 Grouping Layer
To reduce the number of image tokens sent to the LLMs while minimizing the performance loss. Building upon the design in  [26], we propose to leverage the capabilities of the pre-trained Vision Encoders. We add several learnable semantic tokens $\{\mathbf{Sem}_{i}\}_{i=1}^{N}$ before the transformer layer of ViT and concat it with the image segments $\{\mathbf{Img}_{i}\}_{i=1}^{M}$ after the linear projection, where $N$ and $M$ represents the number of semantic groups and original image segments tokens. … The Grouping Block takes the learned semantic tokens and image segment tokens as input and merges all the segment tokens that are assigned to the same semantic components into a single new image segment, based on similarity in the embedding space. To be more specific, we compute the similarity matrix A between the semantic tokens $\{\mathbf{Sem}_{i}\}_{i=1}^{N}$ and image segments tokens $\{\mathbf{Img}_{i}\}_{i=1}^{M}$ through a Gumbel-Softmax operation computed over semantic tokens as
@@ 3.4 Instruction-aware Visual Token Grouping
Stage 2: Visual Instruction Tuning
Once the image tokenizer has been pre-trained and the visual features are aligned with the language model, we move on to the second stage, which focuses on fine-tuning the model for instruction-aware visual token grouping. At this point, we incorporate the grouping block $G$ along with semantic tokens into the Vision Transformer architecture to enable effective visual token grouping based on specific instructions.
@@ 4.1 Datasets
For fair comparison, we conduct experiments on the same datasets as introduced in  [15], which is $\sim$558K image-text pair for visual connector pre-training and $\sim$665K instruct-following data for visual instruct-tuning. Since many image links in the dataset of the instruction tuning stage have expired, compared to the original setting ($\sim$665K), only $\sim$624K data are available. The performance of LLaVA1.5 reported in our analysis is reproduced by ourselves to ensure a fair comparison under the same experimental environment and dataset setting. For downstream tasks, we evaluate our model on GQA [10], TextVQA [19], POPE [12], MMBench [16], ScienceQA [17], MME [8]

### v2

@@ 2.2 Efficient inference for LLM/MLLM
The auto-regressive nature of LLMs poses a great challenge to the deployment of LLMs. The quadratic complexity of computing the attention makes the generation process much slower when the input token is longer. StreamingLLM [27] and FastGen [11] prune the redundant attention computation to simplify the computation. Despite their success, they are designed for the single-modal LLM and have not proven successful when it comes to scenarios involving tokens from other modalities. For improving the efficiency of MLLMs, various works [13, 36, 34] adopt Q-Former to map the images to fixed-length tokens. In the meantime, many works try to train smaller MLLMs with smaller backbone [32, 35] to handle the scenario with less computational resources, MoE-LLaVA [16] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance. Another line of research tries to reduce the number of visual tokens while keeping the backbone of LLM unchanged. Since LLMs contribute most to the computation, the number of input tokens to the LLM becomes the bottleneck of the inference cost. Various works [4, 6] have put effort into increasing the throughput for the vision encoder. DeCo [30] employs 2D adaptive average pooling to down-sample the visual tokens at the patch level. Honeybee [5] proposes to use ResNet Block and deformable attention to conduct the abstraction of the vision tokens. VoCo-LLaMA [31] compress the vision tokens using the LLMs and leverage the attention distillation to let LLMs restore information from the specially defined VoCo tokens instead of the whole image tokens.
@@ 3.1 Recap of Multi-modal Large Language Models
MLLMs aim to develop a powerful model capable of generating responses that follow the instructions given for multi-modal inputs, including visual and textual data. MLLMs are typically composed of three core components: 1) Visual Encoder $E_{v}$: it converts an input image $I\in\mathbb{R}^{H\times W\times 3}$ into a set of distinctive visual embeddings $I_{v}\in\mathbb{R}^{N\times C}$. CLIP-ViT-L/14 with patch size 14 are widely adopted as the visual encoder due to their language-aligned pretrained nature, $N=HW/P^{2}$ denotes the number of visual embeddings. 2) Visual Projector $E_{p}$, parametrized by $W$, which is typically a multi-layer perceptron: this component translates visual embedding $I_{v}$ into the visual token $T_{v}$ in the textual embedding space $T$ with an appropriate dimension for the subsequent language model. This part substantially serves as a tokenizer for the input image 3) Large Lanugae Model $E_{l}$, parameterized by $\phi$: it takes in both visual token $T_{v}$ and textual token $T_{t}$, and produces an appropriate response auto-regressively. For a sequence of responses with length $L$, we compute the probability of the target answers $T_{a}$ by
@@ 3.2 Grouping Layer
To reduce the number of image tokens sent to the LLMs while minimizing the performance loss. Building upon the design in  [28], we propose to leverage the capabilities of the pre-trained Vision Encoders. We add several learnable semantic tokens $\{\mathbf{Sem}_{i}\}_{i=1}^{N}$ before the transformer layer of ViT and concat it with the image segments $\{\mathbf{Img}_{i}\}_{i=1}^{M}$ after the linear projection, where $N$ and $M$ represents the number of semantic groups and original image segments tokens. The Grouping Block takes the learned semantic tokens and image segment tokens as input and merges all the segment tokens that are assigned to the same semantic components into a single new image segment, based on similarity in the embedding space. To be more specific, we compute the similarity matrix A between the semantic tokens $\{\mathbf{Sem}_{i}\}_{i=1}^{N}$ and image segments tokens $\{\mathbf{Img}_{i}\}_{i=1}^{M}$ through a Gumbel-Softmax operation computed over semantic tokens as
@@ 3.4 Instruction-aware Visual Token Grouping
Stage 2: Visual Instruction Tuning Once the image tokenizer has been pre-trained and the visual features are aligned with the language model, we move on to the second stage, which focuses on fine-tuning the model for instruction-aware visual token grouping. At this point, we incorporate the grouping block $G$ along with semantic tokens into the Vision Transformer architecture to enable effective visual token grouping based on specific instructions.
@@ 4.1 Datasets
For fair comparison, we conduct experiments on the same datasets as introduced in  [17], which is $\sim$558K image-text pair for visual connector pre-training and $\sim$665K instruct-following data for visual instruct-tuning. Since many image links in the dataset of the instruction tuning stage have expired, compared to the original setting ($\sim$665K), only $\sim$624K data are available. The performance of LLaVA1.5 reported in our analysis is reproduced by ourselves to ensure a fair comparison under the same experimental environment and dataset setting. For downstream tasks, we evaluate our model on GQA [12], TextVQA [21], POPE [14], MMBench [18], ScienceQA [19], MME [10].